OpenStudy (christopherotten):

A certain forest covers an area of 2,000 square kilometers. Suppose that each year this area decreases by 6%. What is the function that best represents the area of the forest each year and how much area remains after 12 years? Round your answer to the nearest square kilometer. Hint: Use the formula, f(x) = P(1 + r)x. f(x) = 2,000(0.94)x, 952 square kilometers f(x) = 2,000(1.06)x, 4,204 square kilometers f(x) = 2,000(0.06)x, 1,239 square kilometers f(x) = 2,000(0.94)x, 1,432 square kilometers help

OpenStudy (christopherotten):

can I have help I do not understand

OpenStudy (erak):

I would go with A. f(x) = P(1 + r)^x f(x) = 2000(1 - .06)^12 f(x) = 952 km^2